# Identity Preserving 3D Head Stylization with Multiview Score Distillation

### v2

@@ 1 Introduction
  • We propose rank weighing for score tensors on latent VQ-VAE channels to regularize the LD gradients, which we show achieve better input color and ID preservation in the distillation process. This introduces a new approach to control stylization via distillation.
@@ 3.1 Preliminaries
$\mathbb{E}_{x_{0},\epsilon,t}[\ \|\epsilon-\hat{\epsilon}(x_{t},t)\|^{2}]$ (2)
@@ 3.2 ID-Preserving Stylization with Distillation
Rank weighted score tensors. While LD reduces the smoothness issues, we may come across some artifacted results, especially around the hair, ears, and neck on some prompts (Fig. 5, full-rank). Notice that the artifacts are more focused on incorrect color distribution rather than the style itself. Since diffusion latents must also contain color distribution information [52], we decide to investigate the SVD of the score tensors along the VQ-VAE latent channel (4) dimension. We notice that the first rank contains most of the stylization and surpassing the contributions of the lower three ranks can mitigate undesired tints. Hence, we re-weigh the diagonal singular value matrix S with linearly decaying coefficients W from the largest singular value to the smallest. Specifically,
@@
where $\{x_{0}^{\pi}\}$ is defined to have all poses $\{0,1,...,N\}$. However, this is not computationally feasible. Instead, as shown in Fig. 2 (b), we approximate $\{x_{0}^{\pi}\}$ with a 2$\times$2 grid, where each element is a $x_{0}^{\pi}$ with different render pose $\pi$. This way, denoising UNet can correlate between different renders of $\theta$, improving stylization consistency across views. We further employ a depth-conditioned ControlNet [63] to ensure the estimated score does not collapse to a single fused image. We also leverage CFG [20] on each backpropagation path in Fig. 2, but omit from figure for simplicity.
@@ 5 Conclusion
Limitations and future work. A limitation, similar to other methods, is that stylization may not fully preserve earrings and hair accessories, as illustrated in Fig. 12. Another one is the output resolution, which is bounded by the utilized generative model. However, they can be mitigated by utilizing a heavier model. For the future work, we will extend our ID preserving approach and findings about domain adaptation to class-agnostic, large reconstruction models. We believe utilizing LD instead of SDS, score rank-weighing, and grid denoising for such pre-trained models will unlock new research directions in efficient and controllable domain adaptation and stylization.
@@ Appendix A LD objective
Using Eqs. 19 and 16 and change of variables in probability:
$p(x_{0}^{\pi}|y)=p(x_{t}^{\pi}|y)\absolutevalue{\frac{\partial x_{t}}{\partial x_{0}}}^{-1}=\frac{p(x_{t}^{\pi}|y)}{\sqrt{\bar{\alpha_{t}}}}$ (20)
@@
[…]
@@
[…]
7:    $x_{t}^{\pi}\leftarrow\sqrt{\bar{\alpha_{t}}}x_{0}^{\pi}+\sqrt{1-\bar{\alpha_{t}}}n$triangleright $\mathbb{E}_{x_{t}}$
@@
[…]
12:    $\texttt{grad}\leftarrow\mathbf{UW\Sigma V^{\text{T}}}$triangleright rank weighing
@@
[…]
@@
[…]
23:    $\{x_{t}^{\pi}\}\leftarrow\sqrt{\bar{\alpha_{t}}}\{x_{0}^{\pi}\}+\sqrt{1-\bar{\alpha_{t}}}n$triangleright $\mathbb{E}_{\{x_{t}\}}$
@@
[…]
28:    $\texttt{grad}\leftarrow\mathbf{UW\Sigma V^{\text{T}}}$triangleright rank weighing
@@
sample_latent utilizes the mapping network of the generator and maps $z$ to $w^{+}$, later to be fed to the generator. make_grid creates a 2$\times$2 grid with 4 inputs. M is realized with torch.flip(x,dims=[-1]). with no_grad() disables PyTorch’s gradient calculation. Note that each time $x_{0}$ is generated, we implicitly pass it through VQ-VAE to embed it into SD’s latent space.
@@ Appendix D Additional results
Figs. 16, 18, 17, 21, 19 and 20 visualizes the outputs of methods for different prompts in 360-degrees.

### v3

@@ 1 Introduction
  • We propose rank weighing for score tensors on latent VAE channels to regularize the LD gradients, which we show achieve better input color and ID preservation in the distillation process. This introduces a new approach to control stylization via distillation.
@@ 3.1 Preliminaries
$\mathbb{E}_{x_{0},\epsilon,t}[\ \|\epsilon-\hat{\epsilon}(x_{t},t)\|^{2}]\$(2)
@@ 3.2 ID-Preserving Stylization with Distillation
Rank weighted score tensors. While LD reduces the smoothness issues, we may come across some artifacted results, especially around the hair, ears, and neck on some prompts (Fig. 5, full-rank). Notice that the artifacts are more focused on incorrect color distribution rather than the style itself. Since diffusion latents must also contain color distribution information [52], we decide to investigate the SVD of the score tensors along the VAE latent channel (4) dimension. We notice that the first rank contains most of the stylization and surpassing the contributions of the lower three ranks can mitigate undesired tints. Hence, we re-weigh the diagonal singular value matrix S with linearly decaying coefficients W from the largest singular value to the smallest. Specifically,
@@
where $\{x_{0}^{\pi}\}$ is defined to have all poses $\{0,1,...,N\}$. However, this is not computationally feasible. Instead, as shown in Fig. 2 (b), we approximate $\{x_{0}^{\pi}\}$ with a $\times2$ grid, where each element is a $x_{0}^{\pi}$ with different render pose $\pi$. This way, denoising UNet can correlate between different renders of $\theta$, improving stylization consistency across views. We further employ a depth-conditioned ControlNet [63] to ensure the estimated score does not collapse to a single fused image. We also leverage CFG [20] on each backpropagation path in Fig. 2, but omit from figure for simplicity.
@@ 5 Conclusion
Limitations and future work. A limitation, similar to other methods, is that stylization may not fully preserve earrings and hair accessories (Fig. 12). Another one is the output resolution, which is bounded by the utilized generative model. However, they can be mitigated by utilizing a heavier model. For the future work, we will extend our approach of domain adaptation to class-agnostic, large reconstruction models. We believe our findings for such pre-trained models will unlock new research directions.
Acknowledgements. We acknowledge EuroHPC Joint Undertaking for awarding the project ID EHPC-AI-2024A02-031 access to Leonardo at CINECA, Italy.
@@ Appendix A LD objective
Using Eqs. 4 and 1 and change of variables in probability:
[…]
@@
3:  $w^{+}\leftarrow\texttt{sample\_latent($\psi$=0.8)}$
4:  $\pi,\pi^{\prime}\leftarrow\texttt{sample\_pose()}$ $\triangleright$ $\mathbb{E}_{\pi}$
@@
6:  $n,t\leftarrow\texttt{noise\_scheduler(0.70,0.96)}$
7:  $x_{t}^{\pi}\leftarrow\sqrt{\bar{\alpha_{t}}}x_{0}^{\pi}+\sqrt{1-\bar{\alpha_{t}}}n$ $\triangleright$ $\mathbb{E}_{x_{t}}$
@@
11:  $\mathbf{U\Sigma V^{\text{T}}}\leftarrow\mathbf{SVD}\texttt{(grad)}$
12:  $\texttt{grad}\leftarrow\mathbf{UW\Sigma V^{\text{T}}}$ $\triangleright$ rank weighing
@@
18:  $w^{+}\leftarrow\texttt{sample\_latent($\psi$=0.8)}$
19:  $\{\pi\}=\pi^{0},\pi^{1},\pi^{2},\pi^{3}\leftarrow\texttt{sample\_pose()}$ $\triangleright$ $\mathbb{E}_{\{\pi\}}$
@@
22:  $n,t\leftarrow\texttt{noise\_scheduler(0.30,0.80)}$
23:  $\{x_{t}^{\pi}\}\leftarrow\sqrt{\bar{\alpha_{t}}}\{x_{0}^{\pi}\}+\sqrt{1-\bar{\alpha_{t}}}n$ $\triangleright$ $\mathbb{E}_{\{x_{t}\}}$
@@
27:  $\mathbf{U\Sigma V^{\text{T}}}\leftarrow\mathbf{SVD}\texttt{(grad)}$
28:  $\texttt{grad}\leftarrow\mathbf{UW\Sigma V^{\text{T}}}$ $\triangleright$ rank weighing
@@
sample_latent utilizes the mapping network of the generator and maps $z$ to $w^{+}$, later to be fed to the generator. make_grid creates a $\times2$ grid with 4 inputs. M is realized with torch.flip(x,dims=[-1]). with no_grad() disables PyTorch’s gradient calculation. Note that each time $x_{0}$ is generated, we implicitly pass it through VAE to embed it into SD’s latent space.
@@ Appendix D Additional results
Figs. 4, 6, 5, 9, 7 and 8 visualizes the outputs of methods for different prompts in 360-degrees.